PP Products has total assets of $646,819, a total asset turnover rate of 2.97, a debt-equity ratio of 1.61, and a return on equity of 0.28 percent. What is the firm's net income
Answer #1

Debt + equity = Total Assets

Equity = Assets * Equity / [ Equity + Debt ]

= 646819 * 1 / [ 1 + 1.61 ]

= 646819 / 2.61

= 247823

return on equity = Net Income / Equity

0.28 = Net Income / 247823

Net Income = 247823 * 0.28

= 69390.54